# Spectroastrometry and Reverberation Mapping of Active Galactic Nuclei. II. Measuring Geometric Distances and Black Hole Masses of Four Nearby Quasars

### v1

@@ 3 SARM Joint Analysis
The procedure underpinning the SARM joint analysis has been described in a number of previous studies, such as W20, G21b, Li et al. (2022), and Li et al. (2024). What distinguishes our analysis from these earlier studies is that we include the latest improvement of BLR dynamical modeling that takes into account radial-dependent responsivity of the BLR (see Li & Wang 2025 for details). This improvement allows us to accommodate for responsivity weighting in SA and emissivity weighting in RM (e.g., Zhang et al. 2021; Li et al. 2022; Li & Wang 2025). For the sake of completeness, we briefly describe the major points of SARM analysis below.
@@ 3.1 Continuum Light Curve Modeling
We first need to evaluate the continuum fluxes on a sufficiently well sampled time grid, which is required to integrate reprocessed emission from the whole BLR with a wide range of time delays. To this end, we employ the damped random walk (DRW) model to delineate the continuum variability, which is a stationary Gaussian process with well-established procedures to interpolate and extrapolate the observed continuum light curve (e.g., Rybicki & Press 1992; Foreman-Mackey et al. 2017; Li et al. 2018). We use a series of evenly spaced points to represent the continuum reconstruction. This series needs to cover an adequate time period prior to the first epoch of the emission line so as to appropriately model the emission-line flux at any given observation epoch. Specifically, the time length of this period should be equal to or longer than the maximum time delay of the BLR. These continuum points plus two hyper-parameters of the DRW process (the damping time scale and variation amplitude) are treated as free parameters and their best-fit values are determined during the subsequent Bayesian inference (see below).
@@ 3.2 BLR Dynamical Modeling
BLR clouds are distributed within a thick disk, with an opening angle and inclination angle (see Figure 1 of Li et al. 2013). The radial distribution of the clouds follows a shifted gamma distribution, namely, the cloud’s radial distance from the central black hole is given by , where is the Schwarzschild radius, is the mean BLR radius, is the fraction of the inner BLR edge relative to , and is a random number drawn from a gamma distribution with and . The parameter quantifies the standard deviation of the BLR cloud’s radial distance from the central point in units of . It also controls the shape of the radial distribution, which may be made to be exponential (), long-tailed (), or narrow peaked (). Within the opening angle, BLR clouds are distributed uniformly in the azimuthal -direction. In the vertical -direction, the angular dispacement of a cloud is assigned as , where the parameter controls the extent to which clouds are clustered over the outer face of the BLR disk and is a random number drawn from a uniform distribution between 0 and 1 (Pancoast et al. 2014).
@@
From photoionization calculations (Korista & Goad 2004), it is known that the reprocessing coefficient of BLR clouds, denoted by , has a radial dependence. The reprocessing coefficient together with the radial distribution of clouds determine the broad-line emissions. To reduce the degrees of freedom, we let be absorbed into the parameterization of the radial distribution and simply set for all clouds. This means that the radial distribution of BLR clouds includes emissivity weights and should not simply be treated as the number density distribution.
@@ 3.3 Calculating Flux, Profile, and Spectroastrometry of the Emission Line
where is the flux ratio of the emission line relative to the underlying continuum, is coordinate of the baseline, is the emissivity-weighted offset of the BLR projected onto the sky, is the offset of the continuum emission, and is the angular-diameter distance. The offset is determined by summing over all clouds as
@@ 3.4 Bayesian Inference
Given the above continuum light curve and BLR dynamical modeling, we can calculate the modeled SA and RM data and compare these against observations to determine the posterior distributions of model parameters using the Markov Chain Monte Carlo (MCMC) technique. The SA data include the profile and differential phases of the broad Pa or Br line. The corresponding likelihood function is written
@@
The joint likelihood function of SARM analysis is the product of Equations (8) and (9), namely, . The posterior probability distribution of model parameters is then , where is the parameter prior. In Appendix B, we list the full model parameters and their priors. We employ the diffusive nested sampling (DNS) algorithm (Brewer et al. 2011) to explore the parameter space and generate Markov chains, which has been shown to be effective at exploring multimodal distributions and strong correlations between parameters. We have incorporated the above BLR dynamical modeling into our previously developed package BRAINS333The latest code is publicly available at https://github.com/LiyrAstroph/BRAINS, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132. (Li et al. 2013, 2018), which utilizes the DNS code CDNest444The latest code is publicly available at https://github.com/LiyrAstroph/CDNest, while the version used in this work is available at https://doi.org/10.5281/zenodo.10700132.. Both BRAINS and CDNest are written in C Language with the standardized message passing interface and are thereby portable to a wide range of supercomputer clusters. From the generated posterior samples of parameters, we assign the best estimates by the median values and the uncertainties by the 68.3% confidence interval.
@@ 4.1 Overview
When running the MCMC, we find that for PDS 456, the continuum phase offset parameter is strongly degenerate with the angular-diameter distance. The Bayesian modeling always tends to fit the whole differential phases with the continuum phase, leaving quite small phase residuals for the BLR emissions. This in turn results in an unrealistically large angular-diameter distance. We therefore switch off the continuum phase offset parameters for PDS 456. As illustrated below, the distance obtained without including the continuum phase offset turns out to be more reasonable. In addition, we also note that the H light curve, shown in the left bottom panel of Figure 2, displays some fine variation patterns that do not have counterparts in the continuum light curve. It seems that these non-echoing H variations are not solely due to noise but their origin is unclear. The analysis reproduces only the averaged variation trend of the H light curve. As a result, there appear quite a few data points deviating from the best reconstruction.
In the averaged differential phase curves (with the continuum phase subtracted) plotted in the second panel of the top left row of Figures 1-4, 3C 273 and NGC 3783 display an approximately S-shaped signal typical for Keplerian rotation, which is characterized by a crest and trough respectively on either side of the center wavelength of the emission line. In contrast, PDS 456 and Mrk 509 show an asymmetric signal dominated by a trough in the differential phase curve, indicative of non-Keplerian motion. In the present dynamical modeling, the latter can be explained by the presence of a fraction of BLR clouds with inflowing/outflowing motion plus anisotropic emissions.
@@ 4.3 Geometric Distances
Using SARM analysis, W20 derived for 3C 273 and G21b derived for NGC 3783. Considering the associated uncertainties of the inferred distances, these values are consistent with our results. However, we stress that the BLR dynamical modeling in both W20 and G21b adopted a globally linear, uniform responsivity, which always results in the same emissivity- and responsivity-weighted BLR sizes (see also Li & Wang 2025). Our improved SARM analysis naturally accounts for the differences in these two BLR sizes and largely alleviates the increased systematic biases in distance measurements. As mentioned in the preceding section, we find that 3C 273 shows strong radial dependence of the responsivity, while this dependence is weak in NGC 3783 (see Figure 5). These results remarkably coincide with the line widths in the mean and rms spectra of the two objects (see Appendix D).
@@ 5.1 Differences between the H and Pa/Br Lines
The fact that the H line is relatively broader means that current SARM analysis underestimates the geometric distance. If simply scaling results with the above estimated differences in the BLR sizes, the geometric distance will increase by roughly 0.07, 0.19, 0.02, and 0.14 dex for PDS 456, 3C 273, Mrk 509, and NGC 3783, respectively. However, we note that a precise evaluation of the influences to the distance measurements requires comprehensive modeling of the H and Pa/Br lines simultaneously, which is beyond the scope of the present work. Nevertheless, our estimate of the Hubble constant remains within a reasonable range, indicating that the discrepancy arising from different emission lines measured by SA and RM is moderate. In future, this issue can be surmounted by near-infrared RM campaigns observing the same broad emission line as SA or a joint analysis of SA and velocity-resolved RM. In the latter approach, one no longer needs to assume the same region for the H and Pa/Br lines but only let the two lines share the same central SMBH (Li et al. 2022).
@@ 5.2 Results of Mrk 509 from Different Years of RM Data
Mrk 509 was monitored for four years between 2019 and 2022 and the H time delay can be detected in three years, except for 2020 due to the absence of clear variation patterns in the RM data (Li et al. 2024). We choose the 2022 RM data for SARM analysis given that this year marked the largest variability. In Figure 11, we illustrate a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance between years 2019, 2021, and 2022. As observed, the best estimates are consistent with each other within uncertainties, although the BLR size in 2021 shows a slightly larger median value. This is as expected since the H time delays in these three years are in overall agreement.
@@ 6 Conclusions
Although our estimated Hubble constant still carries a large uncertainty, the forthcoming upgraded instrument GRAVITY+ will be able to observe more AGNs (up to a few hundred) with higher-quality differential phase measurements (GRAVITY+ Collaboration et al. 2022). A sizeable RM campaign coordinated with GRAVITY+ interferometric observations will enable the SARM approach to achieve a competitive precision for the Hubble constant. In light of the geometric nature of the measured distance, the SARM approach will ultimately offer a promising cosmic probe, independent of all other cosmic tools.
@@ Acknowledgements
We thank Misty Bentz for kindly sharing the RM light-curve data of NGC 3783 and Hannah Übler for useful comments on the manuscript. YRL acknowledges financial support from the National Key R&D Program of China (2021YFA1600404, 2023YFA1607904), the National Natural Science Foundation of China (12273041), and from the Youth Innovation Promotion Association CAS. JMW acknowledges financial support from the National Natural Science Foundation of China (11991050, 12333003). LCH was supported by the National Key R&D Program of China (2022YFF0503401), the National Science Foundation of China (11991052, 12233001), and the China Manned Space Project (CMS-CSST-2021-A04, CMS-CSST-2021-A06).
We thank the support of the staff of the Lijiang 2.4 m telescope and of the CAHA 2.2m telescope. This work is based on observations collected at the Centro Astronómico Hispanoen Andalucía (CAHA) at Calar Alto, operated jointly by the Andalusian Universities and the Instituto de Astrofísica de Andalucía (CSIC). This work has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
@@ Appendix E The Upper Limit of the Responsivity
In our calculations, the responsivity is parameterized as a power law of radial distance (see Equation 1). The maximum responsivity is limited to according to previous photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014, 2015; Zhang et al. 2021; Li et al. 2024). To test whether this limit affects the parameter inference, we rerun the SARM analysis using an upper limit of and . In Figure 14, we show a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance for these different upper limits. We find that for Mrk 509, PDS 456, and NGC 3783, the results are quite insensitive to the choice of . For 3C 273, the upper limit of tends to yield a smaller BLR size and thereby a smaller angular-diameter distance. Nevertheless, the differences from those of and are about 0.1 dex, still within 1 uncertainties, implying that the influence of on the distance measurement is not significant. A BLR dynamical model with an inclusion of photoionization models will relax the need of present simple parameterization for the responsivity and help further improve SARM analysis. This is beyond the scope of this paper and will be deferred to a future work.

### v2

@@ 3 SARM Joint Analysis
The procedure underpinning the SARM joint analysis has been described in a number of previous studies, such as W20, G21b, Li et al. (2022), and Li et al. (2024). What distinguishes our analysis from these earlier studies is that we include the latest improvement of BLR dynamical modeling that takes into account radial-dependent responsivity of the BLR (see Li & Wang 2025 for details). This improvement allows us to accommodate for responsivity weighting in SA and emissivity weighting in RM (e.g., Zhang et al. 2021; Li et al. 2022; Li & Wang 2025). For the sake of completeness, we briefly describe the major points of the SARM analysis below.
@@ 3.1 Continuum Light Curve Modeling
We first need to evaluate the continuum fluxes on a sufficiently well-sampled time grid, which is required to integrate reprocessed emission from the whole BLR with a wide range of time delays. To this end, we employ the damped random walk (DRW) model to delineate the continuum variability, which is a stationary Gaussian process with well-established procedures to interpolate and extrapolate the observed continuum light curve (e.g., Rybicki & Press 1992; Foreman-Mackey et al. 2017; Li et al. 2018). We use a series of evenly spaced points to represent the continuum reconstruction. This series needs to cover an adequate time period prior to the first epoch of the emission line so as to appropriately model the emission-line flux at any given observation epoch. Specifically, the time length of this period should be equal to or longer than the maximum time delay of the BLR. These continuum points plus two hyperparameters of the DRW process (the damping time scale and variation amplitude) are treated as free parameters, and their best-fit values are determined during the subsequent Bayesian inference (see below).
@@ 3.2 BLR Dynamical Modeling
BLR clouds are distributed within a thick disk, with an opening angle and inclination angle (see Figure 1 of Li et al. 2013). The radial distribution of the clouds follows a shifted gamma distribution, namely, the cloud’s radial distance from the central black hole is given by , where is the Schwarzschild radius, is the mean BLR radius, is the fraction of the inner BLR edge relative to , and is a random number drawn from a gamma distribution with and . The parameter quantifies the standard deviation of the BLR cloud’s radial distance from the central point in units of . It also controls the shape of the radial distribution, which may be made to be exponential (), long-tailed (), or narrow peaked (). Within the opening angle, BLR clouds are distributed uniformly in the azimuthal -direction. In the vertical -direction, the angular displacement of a cloud is assigned as , where the parameter controls the extent to which clouds are clustered over the outer face of the BLR disk and is a random number drawn from a uniform distribution between 0 and 1 (Pancoast et al. 2014).
@@
From photoionization calculations (Korista & Goad 2004), it is known that the reprocessing coefficient of BLR clouds, denoted by , has a radial dependence. The reprocessing coefficient , together with the radial distribution of clouds, determines the broad-line emissions. To reduce the degrees of freedom, we let be absorbed into the parameterization of the radial distribution and simply set for all clouds. This means that the radial distribution of BLR clouds includes emissivity weights and should not simply be treated as the number density distribution.
@@ 3.3 Calculating Flux, Profile, and Spectroastrometry of the Emission Line
where is the flux ratio of the emission line relative to the underlying continuum, is the coordinate of the baseline, is the emissivity-weighted offset of the BLR projected onto the sky, is the offset of the continuum emission, and is the angular-diameter distance. The offset is determined by summing over all clouds as
@@ 3.4 Bayesian Inference
Given the above continuum light curve and BLR dynamical modeling, we can calculate the modeled SA and RM data and compare these against observations to determine the posterior distributions of model parameters using the Markov Chain Monte Carlo (MCMC) technique. The SA data include the profile and differential phases of the broad Pa or Br line. The corresponding likelihood function is written as
@@
The joint likelihood function of SARM analysis is the product of Equations (8) and (9), namely, . The posterior probability distribution of model parameters is then , where is the parameter prior. In Appendix B, we list the full model parameters and their priors. We employ the diffusive nested sampling (DNS) algorithm (Brewer et al. 2011) to explore the parameter space and generate Markov chains, which has been shown to be effective at exploring multimodal distributions and strong correlations between parameters. We have incorporated the above BLR dynamical modeling into our previously developed package BRAINS333The latest code is publicly available at https://github.com/LiyrAstroph/BRAINS, while the version used in this work is available in Li (2024a). (Li et al. 2013, 2018), which utilizes the DNS code CDNest444The latest code is publicly available in https://github.com/LiyrAstroph/CDNest, while the version used in this work is available in Li (2023).. Both BRAINS and CDNest are written in C Language with the standardized message passing interface and are thereby portable to a wide range of supercomputer clusters. From the generated posterior samples of parameters, we assign the best estimates by the median values and the uncertainties by the 68.3% confidence interval.
@@ 4.1 Overview
When running the MCMC, we find that for PDS 456, the continuum phase offset parameter is strongly degenerate with the angular-diameter distance. The Bayesian modeling always tends to fit the whole differential phases with the continuum phase, leaving quite small phase residuals for the BLR emissions. This in turn results in an unrealistically large angular-diameter distance. We therefore switch off the continuum phase offset parameters for PDS 456. As illustrated below, the distance obtained without including the continuum phase offset turns out to be more reasonable. In addition, we also note that the H light curve, shown in the left bottom panel of Figure 2, displays some fine variation patterns that do not have counterparts in the continuum light curve. It seems that these non-echoing H variations are not solely due to noise but their origin is unclear. The analysis reproduces only the averaged variation trend of the H light curve. As a result, there appear a number of data points deviating from the best reconstruction.
In the averaged differential phase curves (with the continuum phase subtracted) plotted in the second panel of the top left row of Figures 1-4, 3C 273 and NGC 3783 display an approximately S-shaped signal typical for Keplerian rotation, which is characterized by a crest and trough respectively on either side of the center wavelength of the emission line. In contrast, PDS 456 and Mrk 509 show an asymmetric signal dominated by a trough in the differential phase curve, indicative of non-Keplerian motion. In the present dynamical modeling, the latter can be explained by the presence of a fraction of BLR clouds with inflowing/outflowing motion plus anisotropic emissions. For PDS 456, there appears an extra red-wing component around 2.24 m, which is not reproduced by our current model. We emphasize that the plotted differential phase curve indeed represents the averaged data from the three baselines with the strongest BLR phase signals (see Appendix C). Notably, this component does not appear across all baselines, possibly implying that it might be caused by some unknown measurement noise or a real BLR component at a special location and velocity that only produces signals along certain baselines. With the available quantity and quality of data we cannot discriminate between these two possibilities. Future higher-quality data will facilitate either a more advanced BLR modeling or a confirmation that we have here a transient observing feature if the extra red-wing component is real.
@@ 4.3 Geometric Distances
Using SARM analysis, W20 derived for 3C 273 and G21b derived for NGC 3783. Considering the associated uncertainties of the inferred distances, these values are consistent with our results. However, we stress that the BLR dynamical modeling in both W20 and G21b adopted a globally linear, uniform responsivity, which always results in the same emissivity- and responsivity-weighted BLR sizes (see also Li & Wang 2025). Our improved SARM analysis naturally accounts for the differences in these two BLR sizes and largely alleviates the increased systematic biases in distance measurements. As mentioned in the preceding section, we find that 3C 273 shows a strong radial dependence of the responsivity, while this dependence is weak in NGC 3783 (see Figure 5). These results remarkably coincide with the line widths in the mean and rms spectra of the two objects (see Appendix D).
@@ 5.1 Differences between the H and Pa/Br Lines
The fact that the H line is relatively broader means that current SARM analysis underestimates the geometric distance. If simply scaling results with the above estimated differences in the BLR sizes, the geometric distance will increase by roughly 0.07, 0.19, 0.02, and 0.14 dex for PDS 456, 3C 273, Mrk 509, and NGC 3783, respectively. However, we note that a precise evaluation of the influences on the distance measurements requires comprehensive modeling of the H and Pa/Br lines simultaneously, which is beyond the scope of the present work. Nevertheless, our estimate of the Hubble constant remains within a reasonable range, indicating that the discrepancy arising from different emission lines measured by SA and RM is moderate. In the future, this issue can be surmounted by near-infrared RM campaigns observing the same broad emission line as SA or a joint analysis of SA and velocity-resolved RM. In the latter approach, one no longer needs to assume the same region for the H and Pa/Br lines but only let the two lines share the same central SMBH (Li et al. 2022).
@@ 5.2 Results of Mrk 509 from Different Years of RM Data
Mrk 509 was monitored for four years between 2019 and 2022 and the H time delay can be detected in three years, except for 2020 due to the absence of clear variation patterns in the RM data (Li et al. 2024). We chose the 2022 RM data for SARM analysis given that this year marked the largest variability. In Figure 11, we illustrate a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance between the years 2019, 2021, and 2022. As observed, the best estimates are consistent with each other within uncertainties, although the BLR size in 2021 shows a slightly larger median value. This is as expected since the H time delays in these three years are in overall agreement.
@@ 6 Conclusions
The currently observed AGNs by GRAVITY remain rare due to the strict magnitude limit (; GRAVITY Collaboration et al. 2017). However, the forthcoming upgraded instrument GRAVITY+ will enable phase reference on a target with using the same observation mode (“on-axis”555Alternatively, with the “off-axis” mode, fainter AGNs () are observable provided there is a bright star () within a 30″ separation serving as the phase reference. However, due to reduced photon fluxes, these AGNs require extended exposure times to achieve an adequate signal-to-noise ratio for BLR differential phases observations.) as the targets reported here. This will permit interferometric observations of a much larger number of AGNs (up to a few hundred; GRAVITY+ Collaboration et al. 2022) and thereby can significantly expand the AGN sample for SARM analysis. The improved sensitivity of GRAVITY+ can also facilitate higher-quality differential phase measurements. Given the feasibility of RM monitoring, those AGNs observable with the on-axis mode are vital for geometric distance determinations. In particular, the Pa line, redshifted into the band for , is generally about 10 times brighter than the Br line for . Therefore, the Pa AGN sample is preferred for future SARM observations with GRAVITY+. In addition, a new visitor interferometry instrument BIFROST at the and bands on VLTI will come into operation within the next few years (Kraus et al. 2022). It is capable of resolving the prominent Pa and Pa lines in low- AGNs (). Moreover, advancements in fringe-tracking (FT) methodologies (such as the hierarchical FT architecture; see Petrov et al. 2024) might enhance the FT limiting magnitude and thus further expand the accessible AGN sample size. Taken together, we would be able to observe sufficient AGN targets for a statistically relevant combination of high-quality SA with RM. The high-quality SA would also allow to eliminate or better model those targets potentially exhibiting unusual features like PDS 456. Although our currently estimated Hubble constant still carries a large uncertainty, a sizable RM campaign coordinated with near-infrared interferometry will enable the SARM approach to achieve a competitive precision for the Hubble constant. In light of the geometric nature of the measured distance, the SARM approach will ultimately offer a promising cosmic probe, independent of all other cosmic tools.
@@ Acknowledgements
We thank Misty Bentz for kindly sharing the RM light-curve data of NGC 3783 and Hannah Übler and the anonymous referee for useful comments on the manuscript. YRL acknowledges financial support from the National Key R&D Program of China (2021YFA1600404, 2023YFA1607904), the National Natural Science Foundation of China (12273041), the China-Chile Joint Research Fund (CCJRF2310), and from the Youth Innovation Promotion Association CAS. JS is supported by the Fundamental Research Funds for the Central Universities, Peking University (7100604896) and the China Manned Space Program (CMS-CSST-2025-A09). JMW acknowledges financial support from the National Natural Science Foundation of China (11991050, 12333003). LCH is supported by the National Key R&D Program of China (2022YFF0503401), the National Science Foundation of China (11991052, 12233001), and the China Manned Space Project (CMS-CSST-2021-A04, CMS-CSST-2021-A06). P.D. is supported by the National Key R&D Program of China (2023YFA1607903) and the China Manned Space Project (CMS-CSST-2025-A07). R.G.P. acknowledges the support of the French Agence Nationale de la Recherche (ANR) through the grant “AGN MELBa” ANR-21-CE31-0011.
We thank the support of the staff of the Lijiang 2.4 m telescope and of the the Centro Astronómico Hispanoen Andalucía (CAHA) 2.2m telescope. This work is based on observations collected at CAHA at Calar Alto, operated jointly by the Andalusian Universities and the Instituto de Astrofísica de Andalucía (CSIC). This work has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.
@@ Appendix E The Upper Limit of the Responsivity
In our calculations, the responsivity is parameterized as a power law of radial distance (see Equation 1). The maximum responsivity is limited to according to previous photoionization calculations (e.g., Korista & Goad 2004; Goad & Korista 2014, 2015; Zhang et al. 2021; Li et al. 2024). To test whether this limit affects the parameter inference, we rerun the SARM analysis using an upper limit of and . In Figure 14, we show a comparison of the obtained BLR size, SMBH mass, and angular-diameter distance for these different upper limits. We find that for Mrk 509, PDS 456, and NGC 3783, the results are quite insensitive to the choice of . For 3C 273, the upper limit of tends to yield a smaller BLR size and thereby a smaller angular-diameter distance. Nevertheless, the differences from those of and are about 0.1 dex, still within 1 uncertainties, implying that the influence of on the distance measurement is not significant. A BLR dynamical model with an inclusion of photoionization models will relax the need for the present simple parameterization for the responsivity and help further improve SARM analysis. This is beyond the scope of this paper and will be deferred to a future work.
Appendix F Comparison with Time Lags from the Cross-correlation Functions
Given a BLR model, we can directly calculate the centroid time lag of the responsivity-weighted transfer function using Equation (A1). In practice, the transfer function usually has a long tail, arising from the responses of BLR clouds at large distances. This long tail causes the centroid time lag of the transfer function to be longer than that from the traditional cross-correlation function (CCF) of the light curves. The latter is conventionally derived by the centroid of the CCF above 80% of the peak value (e.g., Gaskell & Peterson 1987; Peterson et al. 2004). Therefore, to ensure a meaningful comparison, we calculate the time lag from BLR modeling as follows. We firstly generate mock datasets based on the reconstructed light curves from the BLR model by adopting the same measurement errors and sampling cadences as the observed data. We then use the same CCF method to calculate the time lags of these mock datasets.
Figure 15 shows the comparison. The time lags from the two approaches are consistent, except for PDS 456, which shows a relatively shorter time lag from BLR modeling (but still consistent at confidence level). We ascribe this discrepancy to the influence of fine variation patterns in the H light curve of PDS 456 (see Figure 2), which appear to lack counterparts in the continuum variations and thereby cannot be reproduced by RM modeling.